Suppose that the duration of a particular type of criminal trial is known to be normally distributed with a mean of 21 days and a standard deviation of 8 days If one of the trials is randomly chosen, find the probability that it lasted more than 27 days. (Round your answer to four decimal places.) Write the probability statement. P
Answer #1

Solution:

Given that,

\mu = 21

\sigma = 8

a) p ( x > 27 )

= 1 -p ( x <  27 )

= 1 - p (( x - \mu /\sigma ) < ( 27 - 21 / 8 ) )

= 1 - p ( z < 0.75 )

Using z table

=1 - 0.7734

= 0.2266

Probability = 0.2266